# Impact of the Hole Gas on Optically Detected Magnetic Resonance in (Cd,Mn)Te Based Quantum Wells

### v1

@@ V Estimation of effective temperature of magnetic ions Mn2+  and carriers
In order to elucidate the effect of carrier concentration on the ODMR signal obtained on both excitonic features, we exploited possibility to control carrier gas density with additional illumination. The question can be asked whether applying microwave radiation to the sample affects the temperature of the carrier gas, and whether it does not depolarize the carriers by heating them in the typical ODMR measurement. Figure 4(a) shows the spectra of reflectivity in the presence of microwaves ($f$ = 12.9 GHz) in a resonant magnetic field for two different circular polarizations, corrected for the giant Zeeman shift, so as to emphasize the difference in the intensity of the exciton structures. It can be seen that in the $\sigma^{-}$ polarization there are two excitonic features – while in the opposite polarization we do not observe the lines of the charged exciton X+. The lack of reflection of the exciton charged X+   in this polarization also confirms the lack of a differential signal for the corresponding wavelength 4(b). These results prove that even in the presence of microwave radiation, the carrier gas is not heated sufficiently to depolarize its spin in the $B$ = 0.462 T field. This gives us an upper estimate for the temperature of the carrier gas $T_{\mathrm{h}}$. Assuming that the uncertainty of the reflection measurement determines the minimum observable intensity X+   (gray line in the Figure 4(c), we can use the formula (1) to find that the temperature of holes under the influence of microwave radiation $T_{\mathrm{h}}<6$ K, Fig. 4(d). Whether we are dealing with a gas that is fully or only partially polarized obviously also depends on the value of the magnetic field. However, in most of the measurements presented in this paper the smallest magnetic fields of the ODMR measurement had a value of about 0.4 - 0.5 T. Hence, it seems reasonable to assume that in all cases, we are dealing with a polarized hole gas, even when the carrier gas temperature is increased as a result of the sample being under microwave radiation.
@@
As discussed in the Introduction, the most straightforward way to determine the temperature of the Mn subsystem is fitting the Brillouin function to the observed Zeeman shift. However, we could also employ an alternative approach based on the results of the ODMR measurements. This method relies on the observation that the shape of the ODMR spectrum obtained by observing the neutral exciton with lower and higher carrier density is slightly different, but the position of the sharp transition lines is the same. As such they can be therefore used to extract the temperature of the system. Quantitatively, the value of the temperature was extracted from numerical calculations of the whole ODMR spectrum. The figure 5(a) shows the comparison between measured and calculated ODMR spectra for the magnetic field perpendicular to the plane of the QW and two different illumination conditions, which corresponds to different temperatures – it can be seen that they quite well reflect the differences observed in the spectra measured with low and increased carrier gas concentration.
@@
As it was mentioned before, the strict indicator of the temperature is the shape of the ODMR spectrum. As the spectrum itself can be complex, we propose using the measure given by the data central moments. The first central moment (M1) provides information about the center of mass of the spectrum. The third one (M3) is connected to the asymmetry of the spectrum (regarding M1). Both of them are changing with the temperature as the occupation of the manganese levels and then transitions are changing the intensities, Fig. 5(b). The exact values of M1 and M3 are changing with the properties of the sample (strain, composition), however the quantitative behavior with temperature is the same. For each sample and the given set of data it is much easier to calculate the moments and estimate the temperature of the magnetic ions on their basis.
@@
Figure 5: (a) ODMR signal as a shift of neutral exciton, with MW frequency $f$ = 12.9 GHz measured without (X) and with increased carrier density (X’), helium bath temperature 1.6 K; numerical calculation of ODMR spectra for $D$ = 780 neV, $C$ =0 and Mn ion temperature $T$ = 2.68 K i $T$ = 7.5 K; (b) black connected points - change of first (M1) and third (M3) central moment in the function of Mn ion temperature calculated from numerical calculated ODMR spectra, blue and red lines represent M1 and M3 calculated for data measured for low (blue) and increased (red) carrier density and corresponding temperatures of Mn ions. (c) Center of the resonance for neutral exciton measured without (black spheres) and with increased carrier density (blue spheres) for different angle of external magnetic field, colored curves mark the numerically calculated evolution corresponding to different temperatures of Mn ions.
@@ VI Dependence on magnetic field angle
To further test our interpretation, we also performed measurements with varying angles of the external magnetic field. The shape of the ODMR spectrum obtained on the neutral exciton line was examined depending on the angle of the external magnetic field $\theta$ with respect to growth axis. The zero angle corresponds to the measurement in the Faraday configuration (magnetic field perpendicular to the QW plane). The results obtained for the situation with low (X) and increased carrier density (X’) are presented in the figure 5(c). Due to the increased carrier density, a significant change in the thermal distribution of manganese ion states leads to changes in the shape of the ODMR signal and an effective shift of their center. The dependence of the resonance field for two different temperatures, 2.68 K and 7.5 K, determined on the basis of numerically calculated ODMR spectra, was plotted on the graph 5(b) obtaining a very good agreement with the experimental points. Such changes in the resonance field depending on the $\theta$ angle confirm that the temperatures of the ion system in the low and increased carrier density conditions are significantly different. Measurement of the ODMR spectrum and its center as a function of a magnetic field angle confirms the accuracy of the used model and the high temperature of the magnetic ions under MW radiation and with increased carrier density.
@@
Figure 6: (a) Dependence of the ODMR signal for microwave frequency 12.9 GHz observed by examining the position of the neutral (X) and charged (CX) exciton lines on the magnetic field for different field angles relative to the UW1467 sample plane – and simulated spectra ODMR for parameters: $T$ = 7.5 K, $C$ = 0 (X) and $C$ = -3240 neV (CX) , $D$ = 780 neV. The constant $C$ is the carrier-ion interaction constant. (b) Positions of the ODMR signal maxima for the charged and neutral exciton depending on the angle of the magnetic field relative to the sample plane (0 means Faraday configuration, 90 – Voigt); (c) The distance of the resonance magnetic fields measured by observing the lines X   and CX   (Knight shift) depending on the angle of the magnetic field; (d) ODMR signal width measured by observing X   and CX, black curve represents simulated values.
@@
Figure 6(a) shows the ODMR signals depending on the magnetic field and for different angles of the field relative to the plane of the sample for charged exciton. Both the ODMR signal obtained by observing the X   and CX   lines change the location of the center of gravity depending on the direction of the magnetic field, fig. 6(b).
@@
For the interactions with the carriers (hole gas), a simple Heisenberg model was used, adding to the Hamiltonian mean field exchange term:
$H_{p-d}=-\beta n|\varphi(z)|^{2}\langle\hat{s}\rangle\cdot\hat{S}_{Mn}=C\langle\hat{s}\rangle\cdot\hat{S}_{Mn}$ (3)
where $\beta$ is exchange integral between hole and Mn ion, $n$ – carrier density, $|\varphi(z)|^{2}$ distribution of carriers along growth axis $z$. Here, we make a justified assumption that the magnetic ions in the QW are localized and interacting with the local density of delocalized carriers – $n|\varphi(z)|^{2}$. It is noteworthy that in this model we assume that Mn2+ ion is coupled with the fully polarized hole gas of a certain density, not to the single, localized carrier. Besides the aforementioned polarizability measurements, this assumption is also justified because we do not see any ODMR signal shifted towards higher magnetic fields. Such signal could be a result of the interaction with the hole (i.e. localized one) anti-parallel to the ground spin state. It is not our case, therefore, we use mean-field approximation, and $\langle\hat{s}\rangle$ denotes the average over the whole gas of holes. The shift towards lower magnetic fields is connected only to the effective magnetic field derived from polarized holes.
@@
In the case of a magnetic field perpendicular to the plane of the well ($\theta=0^{\circ}$), the shift of the resonances measured between both exciton complexes is about 23 mT – which is reproduced for the constant $C$ = -3240  neV. This value, according to the equation (3), corresponds to a hole gas concentration of 4.05 $\cdot$1010 cm-2. From the fitting of excitonic lines of neutral and charged state in zero-field reflectance spectrum we estimated based on Kossacki et al., 1999 and Kossacki et al., 2004 the carrier concentration to be about 3 $\cdot 10^{10}$ cm-2 (with additional illumination). This value obtained from the Knight shift is slightly higher than estimated from reflection spectra. Assuming that the carrier density obtained from reflectance spectrum is a mean over the laser spot area and at the same time, the Knight shift is a local result for areas with the excess of carriers we can estimate what is the percentage of those areas – it is about 75%.
@@ VII Discussion
Even for such a simple toy model, the agreement with the experimental data is satisfactory. One of the possible reasons for the remaining deviations is the strong assumption of the equal temperature of the magnetic ions probed by X and CX, or the parameters remain constant with changing magnetic field angle. The exchange integral in this system is isotropic. However, the $C$ parameter contains the local density of the carrier wavefunction, which, in contrast, can depend on the $\theta$ angle. We consider a uniform distribution of charges in the plane of the well. At the same time, the literature data Golnik et al. (2004); Maślana et al. (2006); Łopion et al. (2020) and the fact of simultaneous observation of different ODMR signals on the X   and CX   lines show that this is not true. Most likely, in the plane of the well, there are areas with an excess of carriers, from which the optical signal comes from the recombination of charged excitons, and areas with a lower carrier density. In this case, the size of the areas is submicron and in the measurement of photoluminescence or reflection in the macro scale it is impossible to separate them.
The observation about the spatial separation of X and CX   goes along with the long-running discussion on that topic. Significant works, such as Moody et al., 2014; Brunhes et al., 1999, have confirmed the correlations between charged and neutral excitons, which is serious evidence for the lack of (or significantly small) separation between the two. However, the level of the separation or its absence can be connected to the (in-) homogeneity of the Coulomb potential in the QW layer. This Coulomb disorder may be a result of the inhomogeneity of the surface of the sample for the systems with low separation between QW and the surface Łopion et al. (2020); Maślana et al. (2006) or inhomogeneity of the QW layer itself. In our case we believe it is a first case scenario – as we observe the possibility of change of the Coulomb disorder with changing the additional illumination and screening the surface potential.
In the next part we discuss the temperatures of the subsystems having a part in the spin relaxation processes observed with the ODMR signal.
@@
For increased carrier density, the situation becomes more complex. In such a case the two different transitions visible the reflectance spectrum – neutral (X) and positively charged exciton (CX) give different characteristics of ODMR signal. In this case, we assume that the optical signal from charged and neutral exciton comes from spatially separated (in plane of QW) regions differing in the carrier density. In the regions with the higher density of carriers, where the optical signal of charged exciton comes from, the ions experience an effective exchange field of polarized hole gas, inducing ODMR signals shifting towards lower external magnetic field. However, even the ions probing the "neutral" regions experience interactions with the increased density of carriers. In this case, the change of the ODMR spectrum is fine but significant. It becomes even more clear when we observe the behavior of the center of mass of the ODMR spectrum with magnetic field angle. The temperature of the magnetic ions estimated based on the ODMR measurements seems to be higher than the temperature derived from giant Zeeman splitting in the same conditions. This means that these two measurements (ODMR and Zeeman splitting) examine different groups of ions. Part of the ions absorbing radiation microwave has a greater contribution to the ODMR spectrum and, simultaneously, a higher temperature than other ions, contributing mainly to building magnetization in the sample. In general, all ions in the QW should absorb microwave radiation the same way. However, there may be differences in rates of energy and spin transfer. A less effectively relaxing population would be favored in creating the ODMR signal. The others, due to increased interactions between them, would be less active in resonance studies, as they relax much faster and their lines are broadened.
@@
Table 1: Table summarizing the temperatures of different subsystems in the QW measured or estimated with different methods. The temperature of the helium bath in all cases was about 1.6 K and in the case of pulsed microwave excitation it slightly increased by about 5-10%. In order to be able to compare possibly similar measurement conditions in the table the microwave excitation means excitation with $f$ = 12.9 GHz, $P_{\mathrm{MW}}$ = $-10$ dBm. As the power emitted by the antenna and absorbed by the sample may differ for different frequencies, it is difficult to compare carrier or ion temperatures in such cases.
Previous works have shown that different subsystems present in the sample – manganese ions, charge carriers, and crystal lattice –can vary in effective temperature Scherbakov et al. (2001); König et al. (2000); Scherbakov et al. (1999). Here we deal with even more complex situation in which manganese ions do not form homogenous system. A summary of the temperature results obtained in this paper is provided in the table 1.
@@
With increased carrier density, the temperatures concluded from the ODMR measurements – also from the measurements with different magnetic field angles – are higher than the temperatures derived from the giant Zeeman splitting with the same conditions. As mentioned before, presumably, a group of ions in the sample is somehow separated and needs a longer time to relax (reorient itself in the magnetic field) after being excited by a microwave pulse, but also to transfer energy in the form of temperature. This population would thus be more active in magnetic resonance signals. The rest would be less active in resonance studies, as they return to equilibrium much faster and their lines are significantly broadened. Nevertheless, this difference is observed only in the presence of carriers. For increased carrier density the ions which are contributing in ODMR signal thermalize in the ambient of hot carriers, which are excited by MW radiation non-resonantly.

### v2

@@ V Estimation of effective temperature of magnetic ions Mn2+  and carriers
We exploited the possibility of controlling carrier gas density with additional illumination to elucidate the effect of carrier concentration on the ODMR signal obtained on both excitonic features. The question can be asked whether applying microwave radiation to the sample is free of the undesired effects, such as the increase in the effective temperature of the carrier gas and the resulting depolarization of the carriers. Figure 4(a) shows the spectra of reflectivity in the presence of microwaves ($f$ = 12.9 GHz) in a resonant magnetic field for two different circular polarizations, corrected for the giant Zeeman shift, so as to emphasize the difference in the intensity of the exciton structures. In the $\sigma^{-}$ polarization, there are two excitonic features – while in the opposite polarization, we do not observe the lines of the charged exciton X+. To make it more pronounced, we can use the difference of the optical spectra measured in ON and OFF states – such a difference we call differential signal. In a situation without microwaves and in the magnetic fields where the carrier gas is polarized (B>0.1 T), the charged exciton feature is present only in one circular polarization of light. When we apply microwave radiation, the polarization of the hole gas may change, resulting in a non-zero differential signal. However, the lack of reflection feature of the charged exciton X+ in the $\sigma^{+}$ polarization proves that even in the presence of microwave radiation, the carrier gas is not heated sufficiently to depolarize its spin in the $B$ = 0.462 T field. This gives us an upper estimate for the temperature of the carrier gas $T_{\mathrm{h}}$. Assuming that the uncertainty of the reflection measurement determines the minimum observable intensity X+ (gray line in Figure 4(c), we can use the formula (1) to find that the temperature of holes under the influence of microwave radiation $T_{\mathrm{h}}<6$ K, Fig. 4(d). Whether we are dealing with a gas that is fully or only partially polarized also depends on the value of the magnetic field. However, in most of the measurements presented in this paper, the smallest magnetic fields of the ODMR measurement had a value of about 0.4 - 0.5 T. Hence, it seems reasonable to assume that in all cases, we are dealing with a polarized hole gas, even when the carrier gas temperature is increased due to the subjecting of the sample to microwave radiation.
@@
As discussed in the Introduction, the most straightforward way to determine the temperature of the Mn subsystem is fitting the Brillouin function to the observed Zeeman shift. However, we could also employ an alternative approach based on the results of the ODMR measurements. This method relies on the observation that the shape of the ODMR spectrum obtained by observing the neutral exciton with lower and higher carrier density is slightly different. Still, the position of the sharp transition lines is the same. As such they can be therefore used to extract the temperature of the system. Quantitatively, the temperature value was extracted from numerical calculations of the whole ODMR spectrum. Figure 5(a) shows the comparison between measured and calculated ODMR spectra for the magnetic field perpendicular to the plane of the QW and two different illumination conditions, which corresponds to different temperatures – it can be seen that they quite well reflect the differences observed in the spectra measured with low and increased carrier gas concentration. We also observe the emergence of an offset (highlighted in the blue region) corresponding to non-resonant heating when the sample is exposed to MW radiation in the charged regime. This offset is not present in the non-charged regime (black curve). This indicates that the non-resonant heating is primarily driven by the absorption of the microwaves by the carrier gas.
@@
As mentioned before, a strict indicator of the temperature value is the shape of the ODMR spectrum. As the spectrum can be complex, we propose using the measure given by the data central moments. The first central moment (M1) provides information about the center of mass of the spectrum. The third one (M3) is connected to the asymmetry of the spectrum (with respect to M1). Both change with the temperature as the occupation of the manganese levels, and consequently, the transitions change the intensities, Fig. 5(b). The exact values of M1 and M3 change with the properties of the sample (strain, composition). However, the quantitative behavior with temperature is the same. For each sample and the given data set, it is much easier to calculate the moments and estimate the temperature of the magnetic ions on their basis.
@@
Figure 5: (a) ODMR signal as a shift of neutral exciton, with MW frequency $f$ = 12.9 GHz measured without (X) and with increased carrier density (X’), helium bath temperature 1.6 K; numerical calculation of ODMR spectra for $D$ = 780 neV, $C$ =0 and Mn ion temperature $T$ = 2.68 K i $T$ = 7.5 K. In the figure, the offset corresponding to non-resonant heating is marked. To calculate the central moments, it is crucial to subtract this part of the spectrum. (b) black connected points - change of first (M1) and third (M3) central moment in the function of Mn ion temperature calculated from numerical calculated ODMR spectra, thick blue and red lines represent M1 and M3 calculated for data measured for low (blue) and increased (red) carrier density and corresponding temperatures of Mn ions, marked with the dashed lines. (c) Center of the resonance for neutral exciton measured without (black spheres) and with increased carrier density (blue spheres) for different angles of the external magnetic field, colored curves mark the numerically calculated evolution corresponding to different temperatures of Mn ions.
@@ VI Dependence on magnetic field angle
To further test our interpretation, we also performed measurements with varying angles of the external magnetic field. The shape of the ODMR spectrum obtained on the neutral exciton line was examined depending on the angle of the external magnetic field $\theta$ with respect to the growth axis. The zero angle corresponds to the measurement in the Faraday configuration (magnetic field perpendicular to the QW plane). The results obtained for the situation with low (X) and increased carrier density (X’) are presented in the figure 5(c). Due to the increased carrier density, a significant change in the thermal distribution of manganese ion states leads to changes in the shape of the ODMR signal and an effective shift of their center. The dependence of the resonance field for two different temperatures, 2.68 K and 7.5 K, determined based on numerically calculated ODMR spectra, was plotted on the graph 5(b) obtaining a very good agreement with the experimental points. Such changes in the resonance field depending on the $\theta$ angle confirm that the temperatures of the ion system in the low and increased carrier density conditions are significantly different. Measurement of the ODMR spectrum and its center as a function of a magnetic field angle confirms the accuracy of the used model and the high temperature of the magnetic ions under MW radiation and with increased carrier density.
@@
Figure 6: (a) Dependence of the ODMR signal for microwave frequency 12.9 GHz observed by examining the position of the neutral (X) and charged (X+) exciton lines on the magnetic field for different field angles relative to the samples plane – and simulated spectra ODMR for parameters: $T$ = 7.5 K, $C$ = 0 (X) and $C$ = -3240 neV (X+) , $D$ = 780 neV. The constant $C$ is the carrier-ion interaction constant. (b) Positions of the ODMR signal maxima for the charged and neutral exciton depending on the angle of the magnetic field relative to the sample plane (0 means Faraday configuration, 90 – Voigt); (c) The distance of the resonance magnetic fields measured by observing the lines X   and X+   (Knight shift) depending on the angle of the magnetic field; (d) ODMR signal width measured by observing X   and X+, black curve represents simulated values.
@@
Here, we show more differences in the ODMR signals obtained as a change of the excitonic line of neutral and charged exciton. Figure 6(a) shows the ODMR signals depending on the magnetic field and for different angles of the field relative to the plane of the sample. We can observe that the distance between the ODMR signals for the two lines is different depending on the magnetic field angle. Both the ODMR signal obtained by observing the X and X+ lines change the location of the center of gravity with the change of the direction of the magnetic field, Fig. 6(b), and their mutual distance is decreasing.
This was another observation that suggests a difference in the properties of magnetic ions probed with the neutral excitonic feature to those probed with the charged excitonic line. In the next part, we show that the characteristics of the ODMR signal obtained with the charged exciton line can be connected directly to the interactions between probed magnetic ions and charge carriers.
@@
For the interactions with the carriers (hole gas), a simple Heisenberg model was used, adding to the Hamiltonian mean-field exchange term:
$H_{p-d}=-\beta n|\varphi(z)|^{2}\langle\hat{s}\rangle\cdot\hat{\mathrm{S}}_{Mn}=C\langle\hat{s}\rangle\cdot\hat{\mathrm{S}}_{Mn}$ (3)
where $\beta$ is exchange integral between hole and Mn ion, $n$ – carrier density, $|\varphi(z)|^{2}$ distribution of carriers along growth axis $z$.
In this model, we assume that localized magnetic ions in the QW interact with fully polarized hole gas. Two facts justify this assumption. First, from the hole polarizability measurements, we observe that even if the carrier gas temperature is elevated after the absorption of the MW, the carrier gas stays in the fully polarized regime. Additionally, we do not detect any ODMR signal shift towards higher magnetic fields. Such a signal could be observed in the case of interaction with a hole spin anti-parallel to the ground spin state. In the fully polarized hole gas regime, the holes of the different (e.g., antiparallel) spin could be linked to strong localization and coupling with manganese ions. However, we have not observed such a signal in the experiment. That also confirms the high quality of the sample and low density of the localizing sites for carriers.
Therefore, we use mean-field approximation, and the carrier’s spin – $\langle\hat{s}\rangle$ denotes the average over the whole gas of holes. The shift towards lower magnetic fields is connected only to the effective magnetic field derived from polarized holes.
In the case of a magnetic field perpendicular to the plane of the well ($\theta=0^{\circ}$), the shift of the resonances measured between both exciton complexes is about 23 mT – which is reproduced for the constant $C$ = -3240  neV. This value, according to the equation (3), corresponds to a hole gas concentration of 4.05 $\cdot$1010 cm-2.
@@
It was already shown for similar structures, in Kossacki et al., 1999, 2004 that the mutual intensity of the features corresponding to the neutral and charged exciton gives information about the carrier density. On that basis, we estimated the carrier concentration to be about 3 $\cdot 10^{10}$ cm-2 (with additional illumination). The value of carrier density obtained from the Knight shift is slightly higher than estimated from reflection spectra. Assuming that the carrier density obtained from the reflectance spectrum is a mean over the laser spot area and that the Knight shift is a local result for areas with the excess of carriers, we can estimate the percentage of those areas – about 75%.
The change of the center of gravity for ODMR spectra obtained as a change of neutral and charged exciton is shown in Fig 6(c) along with the curve obtained as the difference between the modeled behavior of each ODMR spectrum shown in Fig 6(b) with red and black curves. In Figure 6(d), the changes in the width of the ODMR signal are also shown. Both of those characteristics – the center of gravity and the width are derived from rather complex ODMR spectra. However, as was discussed in detail in the previous part for the neutral regime – the central moments of spectra give a fingerprint of the samples’ condition, i.e. the temperature of magnetic ions. Here again, the first moment M1 (center of gravity) and the second M2 (width) show the agreement between the experimentally obtained data and the modeled spectrum behavior in the magnetic field.
@@ VII Discussion
Even for such a simple toy model, the agreement with the experimental data is satisfactory. One of the possible reasons for the remaining deviations is the strong assumption of the equal temperature of the magnetic ions probed by X and X+, or the parameters remain constant with changing magnetic field angle. The exchange integral in this system is isotropic. However, the $C$ parameter contains the local density of the carrier wavefunction, which, in contrast, can depend on the $\theta$ angle. We consider a uniform distribution of charges in the plane of the well. At the same time, the literature data Golnik et al. (2004); Maślana et al. (2006); Łopion et al. (2020) and the fact of simultaneous observation of different ODMR signals on the X   and X+   lines show that this is not true. Most likely, in the plane of the well, there are areas with an excess of carriers, from which the optical signal comes from the recombination of charged excitons, and areas with a lower carrier density. In this case, the size of the areas is submicron, and in the measurement of photoluminescence or reflection in the macro scale, it is impossible to separate them.
The observation about the spatial separation of X and X+ goes along with the long-running discussion on that topic. Significant works, such as Moody et al., 2014; Brunhes et al., 1999, have confirmed the correlations between charged and neutral excitons, which is serious evidence for the lack of (or significantly small) separation between the two. However, the level of the separation or its absence can be connected to the (in-) homogeneity of the Coulomb potential in the QW layer. This Coulomb disorder may be a result of the inhomogeneity of the surface of the sample for the systems with low separation between QW and the surface Łopion et al. (2020); Maślana et al. (2006) or inhomogeneity of the QW layer itself. In our case we believe it is a first case scenario – as we observe the possibility of change of the Coulomb disorder with changing the additional illumination and screening the surface potential.
In the next part, we discuss the temperatures of the subsystems that contribute to the spin relaxation processes observed with the ODMR signal.
@@
For increased carrier density, the situation becomes more complex. In such a case, the two transitions visible in the reflectance spectrum – neutral (X) and positively charged exciton (X+) give different characteristics of the ODMR signal. In this case, we assume that the optical signal from charged and neutral exciton comes from spatially separated (in plane of QW) regions differing in the carrier density. In the regions with the higher density of carriers, where the optical signal of charged exciton comes from, the ions experience an effective exchange field of polarized hole gas, inducing ODMR signals shifting towards a lower external magnetic field. However, even the ions probing the "neutral" regions experience interactions with the increased density of carriers. In this case, the change in the ODMR spectrum is fine but significant. It becomes even more apparent when we observe the behavior of the center of mass of the ODMR spectrum with a magnetic field angle. The temperature of the magnetic ions estimated based on the ODMR measurements seems to be higher than the temperature derived from giant Zeeman splitting in the same conditions. This means that these two measurements (ODMR and Zeeman splitting) examine different groups of ions. Part of the ions absorbing radiation microwave has a greater contribution to the ODMR spectrum and, simultaneously, a higher temperature than other ions, contributing mainly to building magnetization in the sample. In general, all ions in the QW should absorb microwave radiation the same way. However, there may be differences in rates of energy and spin transfer. A less effectively relaxing population would be favored in creating the ODMR signal. Due to increased interactions between them, the others would be less active in resonance studies, as they relax much faster and their lines are broadened.
@@
Table 1: Table summarizing the temperatures of different subsystems in the QW measured or estimated with different methods. The temperature of the helium bath in all cases was about 1.6 K, and in the case of pulsed microwave excitation, it slightly increased by about 5-10%. In order to be able to compare possibly similar measurement conditions in the table the microwave excitation means excitation with $f$ = 12.9 GHz, $P_{\mathrm{MW}}$ = $-10$ dBm. As the power emitted by the antenna and absorbed by the sample may differ for different frequencies, one should be careful to compare carrier or ion temperatures in such cases.
Previous works have shown that different subsystems present in the sample – manganese ions, charge carriers, and crystal lattice – can vary in effective temperature Scherbakov et al. (2001); König et al. (2000); Scherbakov et al. (1999). Here, we deal with an even more complex situation in which manganese ions do not form a homogenous system. A summary of the temperature results obtained in this paper is provided in the table 1.
@@
With increased carrier density, the temperatures concluded from the ODMR measurements – also from the measurements with different magnetic field angles – are higher than the temperatures derived from the giant Zeeman splitting with the same conditions. As mentioned before, presumably, a group of ions in the sample is somehow separated and needs a longer time to relax (reorient itself in the magnetic field) after being excited by a microwave pulse, but also to transfer energy in the form of temperature. This population would thus be more active in magnetic resonance signals. The rest would be less active in resonance studies, as they return to equilibrium much faster and their lines are significantly broadened. Nevertheless, this difference is observed only in the presence of carriers. For increased carrier density, the ions contributing to the ODMR signal thermalize in the ambient of hot carriers, which are excited by MW radiation non-resonantly.